A coordinator of a degree program in a university assessed a probability distribution for the number of students (x) entering into the program as follows:


X 30 45 60 65 90 100

f(x) 0.10 0.20 0.20 0.30 0.10 0.10


(i) Is the probability distribution valid? Explain


(ii) What is the probability of 60 or fewer students entering into the program?  


1
Expert's answer
2022-04-29T03:13:37-0400

In order to verify that the distribution is valid, we need to check that: "\sum_xf(x)=1". We receive: "0.1+0.2+0.2+0.3+0.1+0.1=1".

(ii) Denote by "X" a random variable that corresponds to that distribution. We need to calculate "P(X\leq60)". We receive: "P(X\leq60)=P(X=30)+P(X=45)+P(X=60)=0.1+0.2+0.2=0.5"

Thus, the probability is "0.5."